Fungicides account for 1/10 of the pesticides used in the U.S. Insecticides account for 1/4 of all the pesticides used in the U.S. The ratio of fungicides to insecticides used in the U.S. can be written 1/10 divided by 1/4. Write this ratio in simplest form.
Answer by fractalier(6550)

Okay 1/10 divided by 1/4 equals
1/10 ÷ 1/4 =
1/10 x 4/1 =
4/10 =
2/5